Hi,
I'm having trouble with the following question:
How many pounds of peanuts worth 70 cents a pound must be mixed with 30 pounds of walnuts worth 90 cents a pound, to produce a mixture which can be sold at 85 cents a pound?
Thanks in advance!

Answer by ewatrrr(24785)

 
Hi
70x + 90(30lb) = 85(x+30lb)
5(30lb)/15 = x, amount of Peanuts